# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v3

@@ 3.1 Heun gravity correspondence
$m_{1}$ $=$ $2{\rm i}M\omega\quad,\quad m_{2}=m_{3}^{*}=-2+2{\rm i}M\omega\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+10{\rm i}M\omega-4\omega^{2}M^{2}$
@@ 4 Wave form and energy flux at 30PN
In this Section we compute the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN, we mean that the polynomials coefficients $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
@@ 4.2 The luminosity
After expanding in harmonics, the corresponding waveform at large distances can be written as
@@
The harmonic coefficients $Z_{\ell m}$ can be written in terms of derivatives of the incoming solution
@@
$b^{0}_{\ell m}={\pi\over 2}\sqrt{(\ell-1)\ell(\ell+1)(\ell+2)}\,Y_{0}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{1}_{\ell m}=\pi\sqrt{(\ell-1)(\ell+2)}Y_{-1}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{2}_{\ell m}=\pi\,Y_{-2}^{\ell m}({\textstyle\frac{\pi}{2}},0),$ (84)
@@
${}_{s}Y_{\ell m}(\theta,\phi)=e^{im\phi}\sin^{2\ell}\left(\frac{\theta}{2}\right)\sqrt{\frac{(2\ell{+}1)(\ell{-}m)!(\ell{+}m)!}{4\pi(\ell{-}s)!(\ell{+}s)!}}\sum_{r=0}^{\ell{-}s}({-})^{\ell{+}m{-}r{-}s}(^{\ell{-}s}_{s})(^{\ell+s}_{r{+}s{-}m})\cot({\textstyle\frac{\theta}{2}})^{2r{+}s{-}m}$ (85)
@@
${d{\cal E}\over dt}=\mu^{2}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}{|Z_{\ell m}|^{2}\over 2\pi\omega^{2}}=\mu^{2}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}\eta_{\ell m}$ (87)
@@
$\left({d{\cal E}\over dt}\right)_{N}=\frac{32\mu^{2}v^{10}}{5M^{2}}=\mu^{2}\,\eta_{22}^{0}$ (88)
with $\eta_{22}^{0}$ the contribution of the leading harmonic mode for $v\to 0$.
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHTK BHPToolkit . In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Figure 2: Luminosity vs velocity. We display the $v$-dependence of the ratio $\eta_{\ell m}/\eta^{0}_{22}$ for the three dominating modes. a: (2,2), b: (3,3), c: (2,1).
The convergence speed of the results is showed in Table 2. The digits in bold style underline the agreement between computations based on the PM hypergeometric representations and MST/numerical results. As expected, the speed of convergence varies with $v$: for $v=0.1$ an agreement on 12 digits is already reached at 10PM order, while for $v=0.4$, contributions at order 30PM are needed to reach a precision on 9 digits.
$k$ $v=0.1$ $v=0.3$ $v=0.4$
5 0.961483706187 0.841473688330 0.887926055076
10 0.961483705502 0.841176615503 0.880582026053
15 0.961483705502 0.841176157786 0.880686334361
20 0.961483705502 0.841176171312 0.880705358744
25 0.961483705502 0.841176171325 0.880705412455
30 0.961483705502 0.841176171325 0.880705411714
@@
Table 2: Convergence of the luminosity results. We display the dependence on the PM order $k$ of the ratio $\eta_{2,2}/\eta^{0}_{2,2}$ computed using the hypergeometric representation. The results are compared against MST/ NUM results displayed in the last line.
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in Fujita:2012cm .
Method $v=0.1$ $v=0.3$ $v=0.4$
@@
HYP 0.961483705502 0.841176171325 0.880705411714
BHTK 0.961483705502 0.841176171325 0.880705411627
Table 3: Luminosity $\eta_{2,2}/\eta^{0}_{2,2}$. Comparison between the different methods.
Method $v=0.1$ $v=0.3$ $v=0.4$
@@
HYP 0.974719282291874 0.950003411484251 1.10991197128
BHTK 0.974719282291999 0.950003411484254 1.10991197120
Table 4: Total luminosity $d{\cal E}/dt/(\mu^{2}\eta^{0}_{2,2})$. Comparison between the different methods.
We notice that in the whole range of $v$ the results obtained with the 30PN hypergeometric representations for the total luminosity $d{\cal E}/dt/(\mu^{2}\eta^{0}_{2,2})$ agree with those obtained by MST and numerical integration up to 11-14 digits.
@@
  • RinRupdRdt.nb reproduces the results of Tables 1, 3 and 4 at order $k=30$. This code imports the data files P30.m, Pt30.m, gg30.m, gk30.m and a30.m, containing the polynomials $P_{0}(y)$, $\widehat{P}_{0}(y)$, the quantities $\gamma$, $\gamma_{\kappa}$ and the $a$ cycle.
  • PNexpansions.nb computes the PN expansions of $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ at any PM order lower than 10. This code imports the data files PNRpp10.m, PNgg10.m, PNgk10.m and PNaa10.m, containing the expansions of the quantities $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ and $a$ at order 10PM.
5 The Heun equation
@@
The dots stand for corrections in $z$. The expansion coefficients of ${\rm Heun}_{\alpha}(z)$ can be obtained recursively, solving the equation order by order in $z$. The Heun differential equation can be related to the quantum SW curve of a ${\cal N}=2$ supersymmetric gauge theory with four fundamental hypermultiplets. To this aim, we write the six Heun parameters in terms of 4 masses $m_{i}$, a gauge coupling $x$ and a Coulomb branch parameter $u$, defined via the identifications777 Equation (89) can be alternatively written in the normal form (5) with $G_{1}(z)=(1-z)^{-\frac{\beta_{1}}{2}}z^{-\frac{\beta_{0}}{2}}(z-\lambda)^{-\frac{\beta_{2}}{2}}\,\Psi(z)$ (92) and $Q_{\rm gauge}$ $=$ $\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}+m_{4}}{2}}\right){}^{2}}{(z-x)^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}-m_{4}}{2}}\right){}^{2}}{z^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{1}+m_{2}}{2}}\right){}^{2}}{(z-1)^{2}}+\frac{m_{3}^{2}+m_{4}^{2}+2m_{1}m_{2}-1}{2(z-1)z}$ (93) $+\frac{(1-x)\left(u+{\textstyle\frac{1}{4}}-\frac{m_{3}^{2}}{2}-\frac{m_{4}^{2}}{2}\right)-{\textstyle\frac{x}{2}}\left(1{-}m_{1}{-}m_{2}\right)\left(1{-}m_{3}{-}m_{4}\right)}{(z-1)z(z-x)}$
@@ 5.2 Connection formulae
Plugging (104) into (5.1) one finds the connection formulae
@@
6 Summary and conclusions
In this paper, we derive a hypergeometric representation of the confluent Heun functions $G_{0}(y)$ or $G_{1}(z)$ that resums the entire $z$ or $y$ dependence of the functions respectively, at any given order in the expansion for small $x$. In the context of gravity, these representations provide exact formulae for the wave form at given order in the PM or PN approximations respectively. The results apply to a general scattering process (along an open or closed trajectory) in the Schwarzschild geometry, where the expansion parameters $z=2M/r$ and $y=2{\rm i}\omega r$ are independent from each other. In a future publication we are going to compare this result with those coming from amplitude computations which, at a given loop order are exact in the velocities. We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}(z)$, $G^{0}_{\alpha}(z)$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by Heun connection matrices. The final picture is summarized in the following scheme
where $B_{\alpha\alpha^{\prime}}$ and $F_{\alpha\alpha^{\prime}}$ are the standard hypergeometric connection matrices relating $z$ to $1/z$ or $1-z$ respectively. The whole non-triviality of the Heun connection formula is codified in the dependence of these matrices on the characteristic function $a(\ell)$. This simple result is a direct consequence of our ansatz for the solutio of the differential equation in terms of a single hypergeometric function and its derivative.
Specifying to circular orbits, we derive the wave form and energy flux at order 30PN. Given that this result is too large to appear in print we also publish ancillary mathematica files with the data needed to perform the computation. When compared against the PN expansions that one can find in literature Blanchet:2013haa ; Fujita:2012cm , these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained using the numerical package BHToolkit, finding an agreement from 9 to 12 digits in the whole range of parameters allowed for bounded orbits.
Similar results can be also obtained for solutions of the non-confluent Heun equation and can therefore be applied to other physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^8] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.
Acknowledgements

### v4

@@ 3.1 Heun gravity correspondence
$m_{1}$ $=$ ${\textstyle\frac{x}{2}}\quad,\quad m_{2}=m_{3}^{*}=-2+{\textstyle\frac{x}{2}}\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+{\textstyle\frac{5x}{2}}+{\textstyle\frac{x^{2}}{4}}$
@@ 4 Waveform and energy flux at 30PN
In this Section we use perturbation theory to compute the probe limit of the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN we mean that the polynomials $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution, are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
@@ 4.2 The luminosity
After expanding in harmonics, the corresponding waveform measured by a far away observer sitting at a point with coordinates $(T,R,\Theta,\Phi)$ can be written as
@@
with $R_{*}$ the tortoise coordinate. The harmonic coefficients $Z_{\ell m}$ can be written in terms of derivatives of the incoming solution
@@
$b^{0}_{\ell m}={\pi\over 2}\sqrt{(\ell^{2}{-}1)\ell(\ell{+}2)}\,Y_{0}^{\ell m}({\textstyle\frac{\pi}{2}},0),~{}~{}b^{1}_{\ell m}=\pi\sqrt{(\ell{-}1)(\ell{+}2)}Y_{-1}^{\ell m}({\textstyle\frac{\pi}{2}},0),~{}~{}b^{2}_{\ell m}=\pi\,Y_{-2}^{\ell m}({\textstyle\frac{\pi}{2}},0),$ (83)
@@
${}_{s}Y_{\ell m}(\theta,\phi)$ $=$ $e^{im\phi}\sin^{2\ell}\left(\frac{\theta}{2}\right)\sqrt{\frac{(2\ell{+}1)(\ell{-}m)!(\ell{+}m)!}{4\pi(\ell{-}s)!(\ell{+}s)!}}$ (84)
$\sum_{r=0}^{\ell{-}s}({-})^{\ell{+}m{+}r{-}s}{{\ell-s}\choose{r}}{{\ell+s}\choose{r+s-m}}\cot({\textstyle\frac{\theta}{2}})^{2r{+}s{-}m}$
@@
${d{\cal E}\over dt}=\mu^{2}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}{|Z_{\ell m}|^{2}\over 2\pi\omega^{2}}=\left({d{\cal E}\over dt}\right)_{N}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}\eta_{\ell m}(v)$ (86)
@@
$\left({d{\cal E}\over dt}\right)_{N}=\frac{32\mu^{2}v^{10}}{5M^{2}}$ (87)
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained with the MST method and numerical integration using the BHTK BHPToolkit code. In all the subsequent tables, the displayed digits are those that are insensitive to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Figure 2: Luminosity vs velocity. We display the $v$-dependence of the ratio between $\eta_{\ell m}$ and the Newton luminosity (87) for the three dominating modes. a: (2,2), b: (3,3), c: (2,1).
The convergence speed of the results for some $(\ell,m)$ modes is showed in Tables 2 and 3. The digits in bold style underline the agreement between computations based on the PM hypergeometric representations and the MST/numerical results. As expected, the speed of convergence varies with $v$: in the case of the $(2,2)$ mode, for $v=0.1$ an agreement at 12 digits is already reached at 10PM order, while for $v=0.4$, the contributions at order 30PM are needed to reach a precision at all the considered digits. A similar behaviour is also found for higher $(\ell,m)$ modes. The last line of each sub-table shows the results coming from the 22PN computation in Fujita:2012cm . We observe that, opposite to what happens if one uses the 22PN expansion, our results agree against those obtained by the MST/Numerical methods for the first 10-12 digits independently of the $\ell$ mode. The difference is evident in the case of large $\ell$, see for example the results for the $(10,1)$ mode in Table 3.
$k$ $\eta_{2,2}(0.1)$ $\eta_{2,2}(0.3)$ $\eta_{2,2}(0.4)$
5 0.961483705513 0.841181503167 0.880796957664
10 0.961483705502 0.841176178745 0.880713162076
15 0.961483705502 0.841176171328 0.880705481610
20 0.961483705502 0.841176171325 0.880705412127
25 0.961483705502 0.841176171325 0.880705411630
30 0.961483705502 0.841176171325 0.880705411627
@@
$k$ $\eta_{2,1}(0.1)$ $\eta_{2,1}(0.3)$ $\eta_{2,1}(0.4)$
5 0.000276130918856 0.00266411244976 0.00572740466492
10 0.000276130918862 0.00266414679182 0.00573006836479
15 0.000276130918862 0.00266414686348 0.00573016433854
20 0.000276130918862 0.00266414686353 0.00573016551089
25 0.000276130918862 0.00266414686353 0.00573016552070
30 0.000276130918862 0.00266414686353 0.00573016552076
BHTK 0.000276130918862 0.00266414686353 0.00573016552076
22PN 0.000276130918862 0.00266414686353 0.00573016504468
$k$ $\eta_{3,3}(0.1)$ $\eta_{3,3}(0.3)$ $\eta_{3,3}(0.4)$
5 0.0127391414224 0.0911202006109 0.166992708659
10 0.0127391414224 0.0911203195830 0.167029390871
15 0.0127391414224 0.0911203195245 0.167029260059
20 0.0127391414224 0.0911203195244 0.167029258239
25 0.0127391414224 0.0911203195244 0.167029258222
30 0.0127391414224 0.0911203195244 0.167029258222
BHTK 0.0127391414224 0.0911203195244 0.167029258222
22PN 0.0127391414224 0.0911203195262 0.167029722596
$k$ $10^{6}\times\eta_{3,2}(0.1)$ $10^{4}\times\eta_{3,2}(0.3)$ $10^{3}\times\eta_{3,2}(0.4)$
5 7.695632933 5.99470001058 2.24151427554
10 7.695632933 5.99470111376 2.24157466375
15 7.695632933 5.99470112089 2.24157691351
20 7.695632933 5.99470112089 2.24157695298
25 7.695632933 5.99470112089 2.24157695337
30 7.695632933 5.99470112089 2.24157695337
BHTK 7.695632933 5.99470112089 2.24157695337
22PN 7.695632933 5.99470112088 2.24157633248
$k$ $10^{6}\times\eta_{3,1}(0.1)$ $10^{6}\times\eta_{3,1}(0.3)$ $10^{5}\times\eta_{3,1}(0.4)$
5 1.1829878889 8.2121836172 1.32887186107
10 1.1829878889 8.2121832569 1.32886658816
15 1.1829878889 8.2121832518 1.32886584761
20 1.1829878889 8.2121832518 1.32886583599
25 1.1829878889 8.2121832518 1.32886583589
30 1.1829878889 8.2121832518 1.32886583589
BHTK 1.1829878889 8.2121832518 1.32886583589
22PN 1.1829878889 8.2121832517 1.32886577283
Table 2: Convergence of the PMk results for $\eta_{\ell,m}$, with $\ell=2,3,1\leq m\leq\ell$, normalized with the Newtonian luminosity. The results obtained using the MST/ NUM methods and the 22PN computation in Fujita:2012cm are displayed in the last two lines of each sub-table.
$k$ $10^{39}\times\eta_{10,1}(0.1)$ $10^{31}\times\eta_{10,1}(0.3)$ $10^{29}\times\eta_{10,1}(0.4)$
5 9.3518204043 8.64616460005 4.13218882098
10 9.3518204044 8.64619838996 4.13187842500
15 9.3518204044 8.64619838996 4.13187842498
20 9.3518204044 8.64619838996 4.13187842498
25 9.3518204044 8.64619838996 4.13187842498
30 9.3518204044 8.64619838996 4.13187842498
BHTK 9.3518204044 8.64619838996 4.13187842498
22PN 9.35182040436 8.64619786463 4.12989060363
Table 3: Convergence of the PMk results for $\eta_{10,1}$ normalized by the Newtonian luminosity. The results obtained using the MST/ NUM methods and the 22PN computation in Fujita:2012cm are displayed in the last two lines.
Finally Table 4 we compare our 30PN results for the energy flux against those obtained by MST/Numerical methods and by the 22PN computation in Fujita:2012cm .
$v=0.1$ $v=0.3$ $v=0.4$
HYP 0.974719282291879 0.950003411484245 1.10991197121
BHTK 0.974719282291999 0.950003411484254 1.10991197120
@@
Table 4: Total luminosity $d{\cal E}/dt$.
We notice that in the whole range of $v$ the results obtained with the 30PN hypergeometric representations for the total luminosity $d{\cal E}/dt$ agree with those obtained by the numerical integration with the MST method up to 11-13 digits.
@@
  • RinRupdRdt.nb: with this file it is possible to compute the $\eta_{\ell m}$ modes importing the data files P30.m, Pt30.m, gg30.m, gk30.m and a30.m, containing the polynomials $P_{0}(y)$, $\widehat{P}_{0}(y)$, the quantities $\gamma$, $\gamma_{\kappa}$ and the $a$ cycle. The results thus obtained will show a slower convergence with respect to the data of Tables 1, 2-3 and 4 which were obtained keeping higher orders than 30PN for $c_{ij},\widehat{c}_{ij}$. The files with these data are quite large and can be obtained by writing to any of the authors.
  • PNexpansions.nb computes the PN expansions of $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ at any PM order lower than 10. This code imports the data files PNRpp10.m, PNgg10.m, PNgk10.m and PNaa10.m, containing the expansions of the quantities $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ and $a$ at order 10PM. If these files are used to reproduce our results above the convergence will be a little slower
5 Summary and conclusions
In this paper, using black hole perturbation theory, we have derived exact formulae for the PMk and PNk expansions (at order $k$ in the soft-limit $\omega M\ll 1$) of the gravitational wave form emitted by a particle moving on an arbitrary trajectory in the Schwarzschild geometry at leading order in the probe limit. The results rely on a novel hypergeometric representation of the solutions of the Heun (or confluent Heun) equation (see Appendix A for the discussion of the non-confluence case). We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}$, $G^{0}_{\alpha}$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by connection matrices. The final picture is summarized in the following scheme
where $B_{\alpha\alpha^{\prime}}$ and $F_{\alpha\alpha^{\prime}}$ are the standard hypergeometric connection matrices relating $z$ to $1/z$ or $1-z$ respectively. The whole non-triviality of the Heun connection formula is codified in the dependence of these matrices on the characteristic function $a(\ell)$. This simple result is a direct consequence of our ansatz for the solution of the differential equation in terms of a single hypergeometric function and its derivative.
Specifying to circular orbits, we derive the probe limit of the wave form and energy flux at order 30PN. In contrast with the PN expansions that one can find in literature Blanchet:2013haa ; Fujita:2012cm , these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. Given that the results are too large to appear in print we collect them in ancillary mathematica files. The results for the waveform have been compared against those obtained using the numerical package BHToolkit, finding an agreement from 9 to 12 digits in the whole range of parameters allowed for bounded orbits.
We remark that our results provide a hypergeometric representations of solutions of the Heun equations of general type, so they apply to any physical problem governed by equations of Heun type. In particular, they can be easily adapted to the study of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes and cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^7] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.
Acknowledgements
Appendix A The Heun equation
@@
The dots stand for corrections in $z$. The expansion coefficients of ${\rm Heun}_{\alpha}(z)$ can be obtained recursively, solving the equation order by order in $z$. The Heun differential equation can be related to the quantum SW curve of a ${\cal N}=2$ supersymmetric gauge theory with four fundamental hypermultiplets. To this aim, we write the six Heun parameters in terms of 4 masses $m_{i}$, a gauge coupling $x$ and a Coulomb branch parameter $u$, defined via the identifications888 Equation (92) can be alternatively written in the normal form (4) with $G_{1}(z)=(1-z)^{-\frac{\beta_{1}}{2}}z^{-\frac{\beta_{0}}{2}}(z-\lambda)^{-\frac{\beta_{2}}{2}}\,\Psi(z)$ (95) and $Q_{\rm gauge}$ $=$ $\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}+m_{4}}{2}}\right){}^{2}}{(z-x)^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}-m_{4}}{2}}\right){}^{2}}{z^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{1}+m_{2}}{2}}\right){}^{2}}{(z-1)^{2}}+\frac{m_{3}^{2}+m_{4}^{2}+2m_{1}m_{2}-1}{2(z-1)z}$ (96) $+\frac{(1-x)\left(u+{\textstyle\frac{1}{4}}-\frac{m_{3}^{2}}{2}-\frac{m_{4}^{2}}{2}\right)-{\textstyle\frac{x}{2}}\left(1{-}m_{1}{-}m_{2}\right)\left(1{-}m_{3}{-}m_{4}\right)}{(z-1)z(z-x)}$
@@ A.2 Connection formulae
Plugging (107) into (A.1) one finds the connection formulae